OpenStudy (anonymous):

How many four-digit numbers can be formed using the digits o,1,2,3,4,5,6,7,8, and 9 if the first digit cannot be 0? Repeated digits are allowed.

OpenStudy (paxpolaris):

9999-999

OpenStudy (anonymous):

I have no idea how to begin. =(

OpenStudy (paxpolaris):

all numbers from 1000 to 9999

OpenStudy (across):

Primero que todo, ¿cuántas posibilidades hay de sacar un numero del 1 al 9?

OpenStudy (anonymous):

[\9 \cdot 9 \cdot 9 \cdot 9\]?

OpenStudy (anonymous):

9 posibilidades

OpenStudy (anonymous):

Me confunde la parte de la repetición!

OpenStudy (across):

Correcto. Ahora, ¿cuántas posibilidades hay de sacar cuatro numeros de una colección de números del 0 al 9?

OpenStudy (mr.math):

I guess it should be \(\left(\begin{matrix}10 \\ 4\end{matrix}\right)-21\).

OpenStudy (paxpolaris):

9thousand 9hundred ninety nine minus 9hundred ninety nine

OpenStudy (anonymous):

9 Posibilidades para el primer número, 9 posibilidades para el segundo, tercero y cuarto.

OpenStudy (anonymous):

Why did you subtracted 21 from that Mr.Math?

OpenStudy (across):

Perdon, quise decir "tres números" ya que el primero fue escogido.

OpenStudy (anonymous):

Entonces, para el primero número tengo 9 posibilidades y para los otros tres tengo 10 posiblidades?

OpenStudy (mr.math):

21 are all the cases at which the first digit is 0.

OpenStudy (across):

¡Exacto!

OpenStudy (mr.math):

Exacto=Exactly, I think.

OpenStudy (anonymous):

Entonces tendría 9*10*10*10?

OpenStudy (across):

Eso es correcto. :)

OpenStudy (anonymous):

Gracias =)

OpenStudy (paxpolaris):

choose doesn't work repeated digits arre allowed

OpenStudy (anonymous):

Pero no se si pueda obtener el mismo resultado con lo que Mr. Math escribió.

OpenStudy (mr.math):

correcto=correct. I speak Spanish! Awesome :D

OpenStudy (mr.math):

So, is my answer correcto or wrongo?

OpenStudy (anonymous):

what's wrong with \( 9 \times 10^3 \) ?

OpenStudy (anonymous):

I'm trying to figure it out Mr.Math. I dont understand the reasoning you're using

OpenStudy (paxpolaris):

9*10*10*10

OpenStudy (anonymous):

This is just fundamental principal of counting ...

OpenStudy (mr.math):

Oh, my answer doesn't really consider all cases. It's wrong!

OpenStudy (anonymous):